# On the information behavior from quadratically coupled accelerated detectors

### v1

@@ I Introduction
It is worth noting that the vast majority of studies involving the UDW detector were done using a linear coupling of the detector with the quantum field. Besides, there are some studies that considered the quadratic coupling hinton1984particle ; takagi1985response ; Takagi1986vacuum , they realized that the detector responds differently from the case coupled linearly. Additionally, the quadratic coupling was considered in various scenarios, namely: fermonic and bosonic fields linearly and quadratically coupled MartinezRenormalized2016 (presenting a renormalization method), entanglement harvesting and divergences in quadratic UDW detector pairs Sachs2017entanglement , scalar and fermionic Unruh Otto engines gray2018scalar , entanglement of two accelerated detectors with an approach different wu2023birth , response of accelerated and nonlinearly coupled detectors sriramkumar2001response and others.
@@
For convenience, we adopt natural units where $c=\hbar=k_{B}=1$ and the metric signature is (+,,,)(+,-,-,-)( + , - , - , - ) throughout the body of this paper.
@@ II.1 The theoretical framework
Before proceeding with the construction of the theoretical framework of this work, it is important to emphasize one last point. As observed by gray2018scalar , for each coupling used we will have a different dimensionality, on the other hand, the same occurs with each coupling constant. In fact, for dimensionless switching functions, it can be shown that $[\lambda_{\phi^{2}}]\equiv-\Delta_{\phi^{2}}=3-d$ (where $d$ is the dimension) considering the quadratic coupling with scalar fields, in units where $[x]=1$. However, it is possible to write a dimensionless constant of the type $\Lambda=\lambda_{\phi^{2}}/\xi^{\Delta_{\phi^{2}}}$, where $\xi$ is a time scale of any perturbative approximation of the detector evolution (for more details see gray2018scalar ). For the case of this work, we have that $d=4$ and the detector evolution time scale is $\xi\sim\Omega^{-1}$, and therefore we obtain, $\Lambda=\lambda_{\phi^{2}}\Omega$.
@@
where the sum over all possible final states allowed us to use the completeness relation $\mathbb{I}=\sum_{\Psi}|\Psi\rangle\langle\Psi|$. Note that in Eq. (5) we use the standard field theoretic technique of normal ordering MartinezRenormalized2016 , where $:\hat{A}:=\hat{A}-\langle 0_{\mathcal{M}}|\hat{A}|0_{\mathcal{M}}\rangle$ for an operator $\hat{A}$. The normal ordering technique is necessary because, when considering quadratic scalar and fermionic detector couplings there are persistent divergences that cannot be regulated either by the switching function or by smearing the detector MartinezRenormalized2016 ; Takagi1986vacuum ; Louko2016 .
@@ II.2 Wightman function
II.3 Probability transition rates
@@
and consequently leading to a geometric series, i.e., $\sum^{\infty}_{k=0}\exp{\left(\frac{-2\pi\Omega k}{a}\right)}=\left[1-\exp{\left(\frac{-2\pi\Omega}{a}\right)}\right]^{-1}$, and we get
@@
where ϕ±()superscriptsubscriptitalic-$\mathcal{R}_{\phi}^{\pm}(\infty)$ is the transition probability rate for an infinite interaction time, for a linear coupling of the scalar field.
@@
$\mathcal{R}^{+}_{\phi^{2}}$$\approx$ $\frac{\Omega\mbox{\rm e}^{2\pi\Omega/a}}{24\pi^{3}}\frac{(\Omega^{2}+a^{2})}{(\mbox{\rm e}^{2\pi\Omega/a}-1)}\Bigg{\{}1+\frac{3}{(\Omega^{2}+a^{2})T^{2}}+$ (34)
@@ III.1.1 The physical model
$\hat{\rho}^{\mathrm{out}}_{\phi^{2}}$ $=$ $\mathcal{\hat{U}}^{(0)}_{\phi^{2}}\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(0)^{\dagger}}_{\phi^{2}}+\mathcal{\hat{U}}^{(1)}_{\phi^{2}}\hat{\rho}_{\mathrm{in}}+\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(0)^{\dagger}}_{\phi^{2}}+\mathcal{\hat{U}}^{(1)}_{\phi^{2}}\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(1)^{\dagger}}_{\phi^{2}}+\mathcal{\hat{U}}^{(2)}_{\phi^{2}}\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(2)^{\dagger}}_{\phi^{2}}+\mathcal{O}(\lambda^{3}_{\phi^{2}}).$ (37)
@@
[…]
@@
and omitting from now on the superscript “$\mathrm{out}$” to simplify notation, finally, one obtains
@@
One highlight that the terms proportional to $\lambda^{2}_{\phi^{2}}$ to have trace null. Here in this setup $\mathbf{Re}(\mathcal{G}^{-})=\mathbf{Re}(\mathcal{G}^{+})$, thus note that,
@@ III.1.2 Quantum coherence
Developing the expression considering $o$ by permuting between the values of $g$ and $e$, we have
@@
Now, using the ratio given by Eq. (52), defining the following dimensionless parameters: $\Lambda=\lambda_{\phi^{2}}\Omega$, $\overline{a}=a/\Omega$, $\sigma=\Omega T$, and ¯$\overline{\mathcal{R}}^{\pm}_{\phi^{2}}=\mathcal{R}^{\pm}_{\phi^{2}}/\Omega$; and using the Eq. (53), one obtains
@@ III.1.4 Numerical results
Aiming at a more direct visual comparison, we plot Fig. 1(e). This plot shows the behavior of coherence for the two types of couplings, both as a function of the parameter $\overline{a}$, where we adopt $\lambda=\Lambda=1.25\times 10^{-2}$. Note that, for these conditions, quantum coherence for the quadratic coupling case decreases rapidly, unlike the linear case, which becomes imperceptible to its degradation. This indicates that the loss of coherence happens much faster when the detector is quadratically coupled with the scalar field. This highlights the difference in strength between the two types of couplings in question.
Figure 1: Behavior of $\mathcal{Q}^{l^{1}}$ as a function of the parameter $\overline{a}$: (a) for quadratic coupling with different values of $\Lambda$, and (b) for linear coupling with different values of $\lambda$. Dependence of $\mathcal{Q}^{l^{1}}$ on (c) $\Lambda$ and (d) $\lambda$ for different values of $\overline{a}$. (e) Plot of $\mathcal{Q}^{l^{1}}_{\phi^{2}}$ and $\mathcal{Q}^{l^{1}}_{\phi}$ as functions of $\overline{a}$. In all cases, the parameters $\theta=\pi/2$ and $\sigma=10$ are held constant.
@@
For a more direct visual comparison, we plot Fig. 2(c). This plot shows the behavior of coherence for the two types of coupling, both as a function of the qubit polar angle, where we adopt $\lambda=\Lambda=1.25\times 10^{-2}$. Note that for these conditions, the quantum coherence for the quadratic coupling case rapidly decreases the oscillation amplitude, unlike the linear case, which becomes imperceptible to its oscillation amplitude degradation.
Figure 2: The representation of $\mathcal{Q}^{l^{1}}$ as a function of the polar angle $\theta$: (a) for quadratic coupling with different values of $\Lambda$ and (b) for linear coupling with different values of $\lambda$. (e) Plot of $\mathcal{Q}^{l^{1}}_{\phi^{2}}$ and $\mathcal{Q}^{l^{1}}_{\phi}$ as functions of the $\theta$. In all cases, the parameters $\overline{a}=100$ and $\sigma=10$ are held constant.
@@
For a more direct visual comparison, we present Fig. 3(e) and Fig. 3(f), which illustrate the behavior of the internal state probabilities for both types of coupling as functions of the parameter $\theta$. In these plots, we adopt $\lambda=\Lambda=1.25\times 10^{-2}$. It is evident that, under these conditions, the probabilities associated with the internal states in the case of quadratic coupling exhibit a more rapid decay in oscillation amplitude. In contrast, the linear coupling case demonstrates a significantly slower degradation, with the reduction in oscillation amplitude being almost imperceptible.
Figure 3: Plot of the probabilities (a) $P^{g}_{\phi^{2}}$ and (b) $P^{e}_{\phi^{2}}$ as functions of the polar angle $\theta$ for different values of $\Lambda$. Plot of the probabilities (c) $P^{g}_{\phi}$ and (d) $P^{e}_{\phi}$ as functions of the polar angle $\theta$ for different values of $\lambda$. Plots of (e) $P^{g}_{\phi}$ and $P^{g}_{\phi^{2}}$, and (f) $P^{e}_{\phi}$ and $P^{e}_{\phi^{2}}$ as functions of $\theta$. In all cases, the parameters $\overline{a}=100$ and $\sigma=10$ are held constant.
@@ III.2.1 The physical model
$\hat{\rho}^{\mathrm{out}}_{\phi^{2}}$ $=$ $\mathcal{\hat{U}}^{(0)}_{\phi^{2}}\hat{\rho}_{\mathrm{in},I}\mathcal{\hat{U}}^{(0)^{\dagger}}_{\phi^{2}}+\mathcal{\hat{U}}^{(1)}_{\phi^{2}}\hat{\rho}_{\mathrm{in},I}+\hat{\rho}_{\mathrm{in},I}\mathcal{\hat{U}}^{(0)^{\dagger}}_{\phi^{2}}+\mathcal{\hat{U}}^{(1)}_{\phi^{2}}\hat{\rho}_{\mathrm{in},I}\mathcal{\hat{U}}^{(1)^{\dagger}}_{\phi^{2}}+\mathcal{\hat{U}}^{(2)}_{\phi^{2}}\hat{\rho}_{\mathrm{in},I}\mathcal{\hat{U}}^{(2)^{\dagger}}_{\phi^{2}}+\mathcal{O}(\lambda^{3}_{\phi^{2}}).$
@@
$\hat{\rho}_{I,\phi^{2}}^{\mathrm{out}}$ $=$ 12[$\frac{1}{2}\left[1+\lambda^{2}_{\phi^{2}}\left(\mathcal{F}^{-}_{\phi^{2}}-2\mathbf{Re}\left(\mathcal{G}^{-}_{\phi^{2}}\right)\right)\right]|g\rangle\langle g|$ (75)
$+$ 12[$\frac{1}{2}\left[1+\lambda^{2}_{\phi^{2}}\left(\mathcal{F}^{+}_{\phi^{2}}-2\mathbf{Re}\left(\mathcal{G}^{+}_{\phi^{2}}\right)\right)\right]|e\rangle\langle e|$
@@ III.2.6 Numeric results
Figure 5: Quantum coherence $\mathcal{Q}^{l^{1}}_{I,\phi}$ (for linear coupling) as a function of $\overline{a}$ and $\alpha$, for several values of $\lambda$. We adopted $\sigma=10$.
@@
Thus, to further clarify this discussion, observe Fig. 6(e), where we represent in the same plot the complementarity relation as a function of $\overline{a}$ for both couplings in question, where we adopt $\Lambda=\lambda=1.25\times 10^{-2}$. Now, it is possible to easily visualize the difference between the degradation of wave-particle duality information when the system reaches high accelerations. Revealing that when considering a detector at high accelerations quadratically coupled with a massless scalar field, the loss of information happens more quickly compared to the case when the detector is coupled linearly with a field of the same type.
Figure 6: The representation of $C_{I}$ as a function of the parameter $\overline{a}$: (a) for quadratic coupling with different values of $\Lambda$, and (b) for linear coupling with different values of $\lambda$. Dependence of $C_{I}$ on (c) $\Lambda$ and (d) $\lambda$ for different values of $\overline{a}$. (e) Plot of $C_{I,\phi^{2}}$ and $C_{I,\phi}$ as functions of $\overline{a}$. In all cases, we adopted $\sigma=10$.

### v2

@@ I Introduction
It is worth noting that the vast majority of studies involving the UDW detector were done using a linear coupling of the detector with the quantum field. Besides, there are some studies that considered the quadratic coupling hinton1984particle ; takagi1985response ; Takagi1986vacuum , they realized that the detector responds differently from the case coupled linearly. Additionally, the quadratic coupling was considered in various scenarios, namely: fermonic and bosonic fields linearly and quadratically coupled MartinezRenormalized2016 (presenting a renormalization method), entanglement harvesting and divergences in quadratic UDW detector pairs Sachs2017entanglement , scalar and fermionic Unruh Otto engines gray2018scalar , entanglement of two accelerated detectors with an approach different wu2023birth , response of accelerated and nonlinearly coupled detectors sriramkumar2001response , enhancement of an UDW battery performance through quadratic coupling mukherjee2024enhancement , and others.
@@
For convenience, we adopt natural units where $c=\hbar=k_{B}=1$ and the metric signature is (+,,,)(+,-,-,-) throughout the body of this paper.
@@ II.1 The theoretical framework
Before proceeding with the construction of the theoretical framework of this work, it is important to emphasize one last point. As observed by gray2018scalar , for each coupling used we will have a different dimensionality, on the other hand, the same occurs with each coupling constant. In fact, for dimensionless switching functions, it can be shown that $\lambda_{\phi^{2}}]\equiv-\Delta_{\phi^{2}}=3-d$ (where $d$ is the dimension) considering the quadratic coupling with scalar fields, in units where $[x]=1$. However, it is possible to write a dimensionless constant of the type $\Lambda=\lambda_{\phi^{2}}/\xi^{\Delta_{\phi^{2}}}$, where $\xi$ is a time scale of any perturbative approximation of the detector evolution (for more details see gray2018scalar ). For the case of this work, we have that $d=4$ and the detector evolution time scale is $\xi\sim\Omega^{-1}$, and therefore we obtain, $\Lambda=\lambda_{\phi^{2}}\Omega$.
@@
where the sum over all possible final states allowed us to use the completeness relation $\mathbb{I}=\sum_{\Psi}|\Psi\rangle\langle\Psi|$. Note that in Eq. (5) we use the standard field theoretic technique of normal ordering MartinezRenormalized2016 , where $\hat{A}:=\hat{A}-\langle 0_{\mathcal{M}}|\hat{A}|0_{\mathcal{M}}\rangle$ for an operator $\hat{A}$. The normal ordering technique is necessary because, when considering quadratic scalar and fermionic detector couplings there are persistent divergences that cannot be regulated either by the switching function or by smearing the detector MartinezRenormalized2016 ; Takagi1986vacuum ; Louko2016 .
@@ II.2 Wightman function
II.3 Transition probability rates
@@
and consequently leading to a geometric series, i.e., $\sum^{\infty}_{k=0}\exp{\left(\frac{2\pi\Omega k}{a}\right)}=\left[1-\exp{\left(\frac{2\pi\Omega}{a}\right)}\right]^{-1}$, and we get
@@
where $\mathcal{R}_{\phi}^{\pm}(\infty)$ is the transition probability rate for an infinite interaction time, for a linear coupling of the scalar field.
@@
$\mathcal{R}^{+}_{\phi^{2}} \displaystyle\approx$ $\frac{\Omega\mbox{\rm e}^{2\pi\Omega/a}}{24\pi^{3}}\frac{(\Omega^{2}+a^{2})}{(\mbox{\rm e}^{2\pi\Omega/a}-1)}\Bigg{\{}1+\frac{3}{(\Omega^{2}+a^{2})T^{2}}+$ (34)
@@ III.1.1 The physical model
$\hat{\rho}^{\mathrm{out}}_{\phi^{2}}$ $=$ $\mathcal{\hat{U}}^{(0)}_{\phi^{2}}\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(0)^{\dagger}}_{\phi^{2}}+\mathcal{\hat{U}}^{(1)}_{\phi^{2}}\hat{\rho}_{\mathrm{in}}+\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(1)^{\dagger}}_{\phi^{2}}+\mathcal{\hat{U}}^{(1)}_{\phi^{2}}\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(1)^{\dagger}}_{\phi^{2}}+\mathcal{\hat{U}}^{(2)}_{\phi^{2}}\hat{\rho}_{\mathrm{in}}+\hat{\rho}_{\mathrm{in}}\mathcal{\hat{U}}^{(2)^{\dagger}}_{\phi^{2}}+\mathcal{O}(\lambda^{3}_{\phi^{2}}).$ (37)
@@
$\mathcal{\hat{U}}^{(1)}_{\phi^{2}}$ $=$ $-i\lambda_{\phi^{2}}\int_{-\infty}^{\infty}d\tau\chi(\tau)\mu(\tau)\phi^{2}[x(\tau)];$ (40)
$\mathcal{\hat{U}}^{(2)}_{\phi^{2}}$ $=$ $-\lambda^{2}_{\phi^{2}}\int^{+\infty}_{-\infty}d\tau\int^{+\tau}_{-\infty}d\tau^{\prime}\chi(\tau)\chi(\tau^{\prime})\mu(\tau)\mu(\tau^{\prime})\phi^{2}[x(\tau)]\phi^{2}[x(\tau^{\prime})],$ (41)
@@
and omitting from now on the superscript $\mathrm{out}$” to simplify notation, finally, one obtains
@@
It is noteworthy that the reduced density matrix Eq. (51) is traceless in terms proportional to $\lambda_{\phi^{2}}^{2}$. Here in this setup $\mathbf{Re}(\mathcal{G}^{-})=\mathbf{Re}(\mathcal{G}^{+})$, thus note that,
@@ III.1.2 Quantum coherence
Developing Eq. (54) considering $i$ and $j$ by permuting between the values of $g$ and $e$, we have
@@
Now, using the relation given by Eq. (52), defining the following dimensionless parameters: $\Lambda=\lambda_{\phi^{2}}\Omega$, $\overline{a}=a/\Omega$, $\sigma=\Omega T$, and $\overline{\mathcal{R}}^{\pm}_{\phi^{2}}=\mathcal{R}^{\pm}_{\phi^{2}}/\Omega$; and using the Eq. (53), one obtains
@@ III.1.4 Numerical results
Aiming at a more direct visual comparison, we plot Fig. 1(e). This plot shows the behavior of coherence for the two types of couplings, both as a function of the parameter $\overline{a}$, where implicitly assuming that $\Omega=1$ we can adopt that $\lambda=\Lambda=1.25\times 10^{-2}$. Note that, for these conditions, quantum coherence for the quadratic coupling case decreases rapidly, unlike the linear case, which becomes imperceptible to its degradation. This indicates that the loss of coherence happens much faster when the detector is quadratically coupled with the scalar field. This highlights the difference in strength between the two types of couplings in question.
Figure 1: Behavior of $\mathcal{Q}^{l^{1}}$ as a function of the parameter $\overline{a}$: (a) for quadratic coupling with different values of $\Lambda$, and (b) for linear coupling with different values of $\lambda$. Dependence of $\mathcal{Q}^{l^{1}}$ on (c) $\Lambda$ and (d) $\lambda$ for different values of $\overline{a}$. (e) The comparison between $\mathcal{Q}^{l^{1}}_{\phi^{2}}$ and $\mathcal{Q}^{l^{1}}_{\phi}$ as functions of $\overline{a}$ where $\Omega=1$ is implicitly assumed. In all cases, the parameters $\theta=\pi/2$ and $\sigma=10$ are held constant.
@@
For a more direct visual comparison, we plot Fig. 2(c). This plot shows the behavior of coherence for the two types of coupling, both as a function of the qubit polar angle, where we adopt $\lambda=\Lambda=1.25\times 10^{-2}$ (implicitly assuming that $\Omega=1$). Note that for these conditions, the quantum coherence for the quadratic coupling case rapidly decreases the oscillation amplitude, unlike the linear case, which becomes imperceptible to its oscillation amplitude degradation.
Figure 2: $\mathcal{Q}^{l^{1}}$ as a function of the polar angle $\theta$: (a) for quadratic coupling with different values of $\Lambda$ and (b) for linear coupling with different values of $\lambda$. (c) Plot of $\mathcal{Q}^{l^{1}}_{\phi^{2}}$ and $\mathcal{Q}^{l^{1}}_{\phi}$ as functions of the $\theta$ (implicitly assuming that $\Omega=1$). The parameters $\overline{a}=100$ and $\sigma=10$ are held constant.
@@
For a more direct visual comparison, we present Fig. 3(e) and Fig. 3(f), which illustrate the behavior of the internal state probabilities for both types of coupling as functions of the parameter $\theta$. In these plots, we adopt $\lambda=\Lambda=1.25\times 10^{-2}$ and for this comparison it is implicitly assumed that $\Omega=1$. It is evident that, under these conditions, the probabilities associated with the internal states in the case of quadratic coupling exhibit a more rapid decay in oscillation amplitude. In contrast, the linear coupling case demonstrates a significantly slower degradation, with the reduction in oscillation amplitude being almost imperceptible.
Figure 3: Probabilities (a) $P^{g}_{\phi^{2}}$ and (b) $P^{e}_{\phi^{2}}$ as functions of the polar angle $\theta$ for different values of $\Lambda$. Plot of the probabilities (c) $P^{g}_{\phi}$ and (d) $P^{e}_{\phi}$ as functions of the polar angle $\theta$ for different values of $\lambda$. Plots of (e) $P^{g}_{\phi}$ and $P^{g}_{\phi^{2}}$, and (f) $P^{e}_{\phi}$ and $P^{e}_{\phi^{2}}$ as functions of $\theta$ where it was implicitly assumed that $\Omega=1$. The parameters $\overline{a}=100$ and $\sigma=10$ are held constant.
@@ III.2.1 The physical model
$\hat{\rho}^{\mathrm{out}}_{\phi^{2}}$ $=$ $\mathcal{\hat{U}}^{(0)}_{\phi^{2}}\hat{\rho}_{\mathrm{in},I}\mathcal{\hat{U}}^{(0)^{\dagger}}_{\phi^{2}}+\mathcal{\hat{U}}^{(1)}_{\phi^{2}}\hat{\rho}_{\mathrm{in},I}+\hat{\rho}_{\mathrm{in},I}\mathcal{\hat{U}}^{(1)^{\dagger}}_{\phi^{2}}+\mathcal{\hat{U}}^{(1)}_{\phi^{2}}\hat{\rho}_{\mathrm{in},I}\mathcal{\hat{U}}^{(1)^{\dagger}}_{\phi^{2}}+\mathcal{\hat{U}}^{(2)}_{\phi^{2}}\hat{\rho}_{\mathrm{in},I}+\hat{\rho}_{\mathrm{in},I}\mathcal{\hat{U}}^{(2)^{\dagger}}_{\phi^{2}}+\mathcal{O}(\lambda^{3}_{\phi^{2}}).$
@@
$\hat{\rho}_{I,\phi^{2}}^{\mathrm{out}}$ $=$ $\frac{1}{2}\left[1+\lambda^{2}_{\phi^{2}}\left(\mathcal{F}^{-}_{\phi^{2}}-2\mathbf{Re}\left(\mathcal{G}^{-}_{\phi^{2}}\right)\right)\right]|g\rangle\langle g|$ (75)
$+$ $\frac{1}{2}\left[1+\lambda^{2}_{\phi^{2}}\left(\mathcal{F}^{+}_{\phi^{2}}-2\mathbf{Re}\left(\mathcal{G}^{+}_{\phi^{2}}\right)\right)\right]|e\rangle\langle e|$
@@ III.2.6 Numeric results
Figure 5: Quantum coherence $\mathcal{Q}^{l^{1}}_{I,\phi}$ (for linear coupling) as a function of $\overline{a}$ and $\alpha$, for several values of $\lambda$. Note that in all plots, we assume $\lambda$ to be 10 times larger than $\Lambda$ to visualize significant effects. We adopted $\sigma=10$.
@@
Thus, to further clarify this discussion, observe Fig. 6(e), where we represent in the same plot the complementarity relation as a function of $\overline{a}$ for both couplings in question, where by implicitly assuming that $\Omega=1$ we can assume that $\Lambda=\lambda=1.25\times 10^{-2}$. Now, it is possible to easily visualize the difference between the degradation of wave-particle duality information when the system reaches high accelerations. Revealing that when considering a detector at high accelerations quadratically coupled with a massless scalar field, the loss of information happens more quickly compared to the case when the detector is coupled linearly with a field of the same type.
Figure 6: The representation of $C_{I}$ as a function of the parameter $\overline{a}$: (a) for quadratic coupling with different values of $\Lambda$, and (b) for linear coupling with different values of $\lambda$. Dependence of $C_{I}$ on (c) $\Lambda$ and (d) $\lambda$ for different values of $\overline{a}$. (e) Plot of $C_{I,\phi^{2}}$ and $C_{I,\phi}$ as functions of $\overline{a}$ with the implicit assumption that $\Omega=1$. In all cases, we adopted $\sigma=10$.